Solve the following system of ODE

x'+x-y = e^2t
1
Expert's answer
2020-10-08T16:47:38-0400

Given Equation,

"x'+x-y=e^{2t}"


"\frac{dx}{dt}+x=y+e^{2t}"


Integrated factor, i.f="e^{\int 1.dt}"

="e^t"

solution is given by,

"x\times i.f.=\int i.f.\times (y+e^{2t})dt+c"


"xe^t=\int(y+e^{2t})e^tdt+c"

"xe^t=ye^t+\frac{e^{3t}}{3}+c"


"x=y+\frac{e^{2t}}{3}+ce^{-t}"